How many grams of HNO3 are in 75ml of a 8.0m HNO3 solution
1
Expert's answer
2021-08-25T06:36:56-0400

There should be an assumption that 8.0 m refers to the molaRity, not molaLity as the mass in this case could not be calculated with only given data (the density of the solution is required). However, using the molarity it can be found. The molar mass of HNO3 is 63.01 g/mol.


"m(HNO_3)=0.075L\times\frac{8.0mol(HNO_3)}{1L}\times\frac{63.01g}{1mol(HNO_3)}=38.g"


Answer: 38. g